one sixth of the smallest of three consecutive even integers is three less than than one tenth the sum of the other two even integers. Find the integers
Answer by tommyt3rd(5050)

n/6=((n+2) + (n+4))/10 - 3
n/6=(2n+6)/10-3
n/6=n/5+3/5-3
3-3/5=n/5-n/6
12/5=n/30
n=6*12=72
They are
{72, 74, 76}